OpenStudy (anonymous):

Evaluate: cube root of 0.343 please show workings

Parth (parthkohli):

Note that\[0.343 = \dfrac{343}{1000}\]So you need to essentially find\[\frac{\sqrt[3]{343}}{\sqrt[3]{1000}}\]